# Generic graded contractions of Lie algebras

### v1

@@ 3 Graded contractions via group cohomology
Recall from Remark 20 that, in the terminology of [39], the first summand in equation (10) corresponds to the classes of contractions with support $S$ that satisfy all surviving higher-order identities. Such classes are described, for $\mathbb{F}=\mathbb{R}$, in [39, Theorem 7.1] in terms of “sign invariants”. In sequence (9), the quotient ${\rm Hom}((\mathcal{C}_{S})_{[2]},\{\pm 1\})$ correspond to the “sign invariants of the first kind” as follows. Fix a set-theoretic section $\xi$ of the projection $\mathcal{Z}_{G}\to\mathcal{C}_{S}$. Given a contraction $\varepsilon$ that satisfies all surviving higher-order identities, i.e., $\varepsilon\in{\rm Hom}(\mathcal{I}_{S},\mathbb{R}^{\times})$, we define the two-cocycle $\alpha_{\varepsilon}\in Z^{2}_{sym}(\mathcal{C}_{S},\{\pm 1\})$ by
@@
where $\mathop{\rm sgn}\nolimits$ denotes the usual sign function, which is the projection ×{±$\mathbb{R}^{\times}\twoheadrightarrow\{\pm 1\}$. The class $[\alpha_{\varepsilon}]\in{\rm Ext}(\mathcal{C}_{S},\{\pm 1\})$ is the image of $[\varepsilon]$ under the homomorphism ${\rm Hom}(\mathcal{I}_{S},\mathbb{R}^{\times})/B^{2}_{S}(\mathbb{R}^{\times})\to{\rm Ext}(\mathcal{C}_{S},\{\pm 1\})$ from sequence (6). The corresponding element $\delta_{\varepsilon}\in{\rm Hom}((\mathcal{C}_{S})_{[2]},\{\pm 1\})$ is given by
@@ 4 Variety of graded contractions
$x_{\{g,h\}}x_{\{g,hk\}}-x_{\{h,k\}}x_{\{g,hk\}}=0\quad\mbox{for all}\quad g,h,k\in G.$ (11)
Let $I\subset\mathbb{F}[\mathbb{A}_{G}]$ be the ideal generated by the left-hand sides of the above equations. This ideal is binomial, meaning it is generated by polynomials with two terms. Binomial ideals have a number of remarkable properties, for instance, the radical and the associated primes of a binomial ideal are again binomial. Eisenbud and Sturmfels [12] provide a detailed study of these and other properties and characterize algebraic sets defined by binomial ideals in terms of their cellular decomposition, see [12, Theorem 4.2]. We will subsequently study the cellular decomposition of the variety $X$ of graded contractions. It should be noted that the ideal $I$ is usually not radical, which can be verified using Singular computer algebra system [7]. For instance, $I$ is radical when $G$ is of order $2$, $3$, $4$ or $5$, but not $6$.
For any subset $S\subset\mathcal{P}_{G}$, we consider the affine subspace $\mathbb{A}_{S}:=\{x\in\mathbb{A}_{G}\colon x_{i}\neq 0~{}\forall i\in S\}\subset\mathbb{A}_{G}$ and the torus ${\rm T}_{S}:=\{x\in\mathbb{A}_{S}\colon x_{i}\neq 0~{}\forall i\in S\}$, which is a principal open set in $\mathbb{A}_{S}$. Since $x\in{\rm T}_{S}$ if and only if $\mathop{\rm supp}x=S$, where $\mathop{\rm supp}x=\{i\in\mathcal{P}_{G}\colon x_{i}\neq 0\}$, we have a partition $\mathbb{A}_{S}=\bigcup_{S^{\prime}\subset S}{\rm T}_{S^{\prime}}$. The tori ${\rm T}_{S}$ are called coordinate cells in [12]. The intersection $X\cap{\rm T}_{S}$ consists precisely of the graded contractions, whose support is $S$; it is nonempty if and only if $S\in\mathcal{S}(G)$. It follows that we have the following partition of $X$:
@@
In particular, $X_{\mathcal{P}_{G}}=X\cap{\rm T}_{\mathcal{P}_{G}}=Z^{2}_{sym}(G,\mathbb{F}^{\times})$ is a nonempty principal open set in $X$. Consider the torus ${\rm T}_{G}:={\rm Fun}(G,\mathbb{F}^{\times})$ and its action on the space $\mathbb{A}_{G}$ defined by
$\alpha\cdot(x_{\{g,h\}})_{\{g,h\}\in\mathcal{P}_{G}}=\big{(}(d\alpha)(g,h)\,x_{\{g,h\}}\big{)}_{\{g,h\}\in\mathcal{P}_{G}},$
where recall that $(d\alpha)(g,h)=\alpha(g)\alpha(h)/\alpha(gh)$. In other words, ${\rm T}_{G}$ acts through the homomorphism $d\colon{\rm T}_{G}\to{\rm T}_{\mathcal{P}_{G}}$, $\alpha\mapsto d\alpha$. This action leaves $X$ and each cell $X_{S}$ invariant. Note that the ${\rm T}_{G}$-orbit of the point $\bar{1}:=(1,\dots,1)$ is $B^{2}(G,\mathbb{F}^{\times})=Z^{2}_{sym}(G,\mathbb{F}^{\times})=X_{\mathcal{P}(G)}$.
@@
The closure of the ${\rm T}_{G}$-orbit of $\bar{1}$, which we denote by $X(\bar{1})$, is the unique irreducible component of $X$ containing $\bar{1}$. Moreover, the points of the orbit are regular in $X(\bar{1})$.
@@
Since ${\rm T}_{G}$ is irreducible, so is $X(\bar{1})$, hence it is contained in an irreducible component of $X$. But it contains an open neighbourhood of $\bar{1}$ in $X$, so it must contain any irreducible component of $X$ passing through $\bar{1}$.
The second assertion is a general property of orbit closures, see discussion in Section 2.1. ∎
@@
Consider the grading group $G=\mathbb{Z}_{2}$. The affine space $A_{\mathbb{Z}_{2}}$ is coordinatized by the tuple $(x_{\{0,0\}},x_{\{1,1\}},x_{\{0,1\}})$ and its coordinate algebra is $\mathbb{C}[x_{\{0,0\}},x_{\{1,1\}},x_{\{0,1\}}]$. The variety of contractions is the zero locus of the polynomials
@@
The computations for the grading group $G=\mathbb{Z}_{3}$ are more complicated than those presented in Example 25. The coordinate algebra is the polynomial algebra in six variables, $\mathbb{C}[x_{\{0,0\}},x_{\{1,1\}},x_{\{2,2\}},x_{\{0,1\}},x_{\{0,2\}},x_{\{1,2\}}]$. The variety $X$ of contractions is the zero locus of the polynomials
@@
$x_{\{g,h\}}x_{\{g,hk\}}-x_{\{h,k\}}x_{\{g,hk\}},\quad x_{\{g,h\}}(x_{\{g,h\}}-1)\quad\mbox{for all}\quad g,h,k\in G.$
@@
All the above computations can be efficiently carried out in Singular. Even for the order-8 grading groups, i.e., the number of varibles of $\mathbb{C}[\mathbb{A}_{G}]$ is $36$, these computations are performed within seconds. In Table 1 we present the numerical results for the number of supports depending on the grading group $G$. We also observe an interesting phenomenon that only a small part of subsets of $\mathcal{P}_{G}$ are actually supports.
@@
$2^{|\mathcal{P}(G)|}$ 8 $2^{6}$ $2^{10}$ $2^{10}$ $2^{15}$ $2^{21}$ $2^{28}$ $2^{36}$ $2^{36}$ $2^{36}$
@@
Orbit closures and higher-order identities. The elements of $X(\bar{1})$ are precisely the continuous contractions as defined in [28, Section 4]. In the case $\mathbb{F}=\mathbb{C}$, $\mathfrak{L}^{\varepsilon}$ is a graded degeneration of a Lie algebra $\mathfrak{L}$ for any $\varepsilon\in X(\bar{1})$, so, in particular, it is a contraction of $\mathfrak{L}$ in the sense of Definition 1. Recall from Section 3 that every relation $r=1$ for $r\in\langle\mathcal{R}_{G}\rangle$ can be uniquely rewritten in the form $r_{1}=r_{2}$, where $r_{1}$ and $r_{2}$ are monomials of the same degree in two disjoint sets of variables. Following [38, 39], we call these relations higher-order identities; they include the defining relations (11). Since the elements of $Z^{2}_{sym}(G,\mathbb{F}^{\times})$ satisfy all higher-order identities, so do the elements of the closure, $X(\bar{1})$. It follows from Theorem 19 and [38, Theorem VI.2] that any element of $X(\bar{1})$ is equivalent via normalization to a generalized Inönü–Wigner contraction.
$(\bar{1}_{S})_{i}:=\begin{cases}1&\mbox{if }i\in S,\\ 0&\mbox{otherwise}.\end{cases}$
@@
$\alpha\cdot(x_{\{g,h\}})_{\{g,h\}\in S}=\big{(}(d\alpha)(g,h)\,x_{\{g,h\}}\big{)}_{\{g,h\}\in S}.$
@@
The orbit of $\bar{1}_{S}$ is the range of the function $f_{S}\colon{\rm Fun}(G,\mathbb{F}^{\times})\to\mathbb{A}_{S}$, $\alpha\mapsto\alpha\cdot\bar{1}_{S}$. Since $\alpha\cdot\bar{1}_{S}=(d\alpha)\bar{1}_{S}$, the range of $f_{S}$ is precisely $B^{2}_{S}(\mathbb{F}^{\times})$ defined in Section 3. For any monomial $r$ in variables indexed by $i\in S$, denote by $\lambda_{r}$ the restriction of the corresponding polynomial function $\mathbb{A}_{S}\to\mathbb{F}$ to the torus ${\rm Fun}(S,\mathbb{F}^{\times})$. First we claim that $\lambda_{r_{1}}|_{B^{2}_{S}(\mathbb{F}^{\times})}=\lambda_{r_{2}}|_{B^{2}_{S}(\mathbb{F}^{\times})}$ if and only if $r_{1}r_{2}^{-1}\in\mathcal{\langle}\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$.
@@
It is clear that $\lambda_{r_{1}}|_{B^{2}_{S}(\mathbb{F}^{\times})}=\lambda_{r_{2}}|_{B^{2}_{S}(\mathbb{F}^{\times})}$ if and only if the function $\lambda_{r_{1}}/\lambda_{r_{2}}$ maps $B^{2}_{S}(\mathbb{F}^{\times})$ to $1$, and the latter condition is equivalent to $r_{1}r_{2}^{-1}\in B^{2}_{S}(\mathbb{F}^{\times})^{\perp}\subset\mathcal{F}_{S}$.
Theorem 19 and Corollary 21 give us that the kernel of the restriction map ${\rm Hom}(\mathcal{Z}_{S},\mathbb{F}^{\times})\to{\rm Hom}(\mathcal{K}_{S},\mathbb{F}^{\times})$ is $B^{2}_{S}(\mathbb{F}^{\times})$. Since $\mathcal{Z}_{S}=\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$ and $\mathcal{K}_{S}=(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})/\langle\mathcal{R}_{S}\rangle$, the kernel of the restriction map ${\rm Hom}(\mathcal{F}_{S},\mathbb{F}^{\times})\to{\rm Hom}(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S},\mathbb{F}^{\times})$ is also $B^{2}_{S}(\mathbb{F}^{\times})$. By definition, this kernel is $(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})^{\perp}$, so we get $B^{2}_{S}(\mathbb{F}^{\times})=(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})^{\perp}$ and hence $B^{2}_{S}(\mathbb{F}^{\times})^{\perp}=((\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})^{\perp})^{\perp}$. Finally, $\mathcal{F}_{S}/(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})\simeq\mathcal{I}_{S}$ is a free abelian group, so its points are separated by homomorphisms to $\mathbb{F}^{\times}$. Therefore, $((\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})^{\perp})^{\perp}=\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$, which completes the proof of the claim.
Let $I$ be the vanishing ideal of the closure of $B^{2}_{S}(\mathbb{F}^{\times})$ in $\mathbb{A}_{S}$. Since $B^{2}_{S}(\mathbb{F}^{\times})$ is dense in its closure, we have
$I=\{\varphi\in\mathbb{F}[\mathbb{A}_{S}]\mid\varphi|_{B^{2}_{S}(\mathbb{F}^{\times})}=0\}.$
A binomial $r_{1}-r_{2}$ arises from a surviving higher-order identity if and only if $r_{1}r_{2}^{-1}\in\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$. It follows from the above claim that all these binomials $r_{1}-r_{2}$ belong to $I$. Modulo a linear combination of such binomials, any $\varphi\in\mathbb{F}[\mathbb{A}_{S}]$ is congruent to a linear combination of monomials $r_{j}$ such that $r_{i}r_{j}^{-1}\notin\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$ for $i\neq j$. But the restrictions $\lambda_{r_{j}}|_{B^{2}_{S}(\mathbb{F}^{\times})}$ are pairwise distinct multiplicative characters of $B^{2}_{S}(\mathbb{F}^{\times})$, hence they are linearly independent over $\mathbb{F}$. Therefore the ideal $I$ is spanned by the binomials $r_{1}-r_{2}$ arising from surviving higher-order identities. This finishes the proof of the first claim of the theorem
@@
The group $\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$ of surviving higher-order identities is a free abelian group of rank $\leqslant|S|$, hence the vanishing ideal of the closure of the orbit of any $\varepsilon\in{\rm T}_{S}$ is generated by the binomials $\varepsilon(r_{2})r_{1}-\varepsilon(r_{1})r_{2}$, where $r_{1}r_{2}^{-1}$ ranges over a generating set of $\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$. Moreover, when the group $G$ is finite, a generating set of $\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$ can be efficiently computed.
@@
The contracted algebra $\mathcal{L}^{\varepsilon}$ is a graded degeneration of $\mathcal{L}$ if and only if $\varepsilon$ satisfies all surviving higher-order identities for $S={\rm Supp}_{2}(\Gamma)$. In particular, if ${\rm Supp}_{2}(\Gamma)=\mathcal{P}_{G}$, then the graded degenerations are precisely the graded contractions in $X(\bar{1})$.

### v2

@@ 3 Graded contractions via group cohomology
Recall from Remark 20 that, in the terminology of [48], the first summand in equation (10) corresponds to the classes of contractions with support $S$ that satisfy all surviving higher-order identities. Such classes are described, for $\mathbb{F}=\mathbb{R}$, in [48, Theorem 7.1] in terms of “sign invariants”. In sequence (9), the quotient ${\rm Hom}((\mathcal{C}_{S})_{[2]},\{\pm 1\})$ corresponds to the “sign invariants of the first kind” as follows. Fix a set-theoretic section $\xi$ of the projection $\mathcal{Z}_{G}\to\mathcal{C}_{S}$. Given a contraction $\varepsilon$ that satisfies all surviving higher-order identities, i.e., $\varepsilon\in{\rm Hom}(\mathcal{I}_{S},\mathbb{R}^{\times})$, we define the two-cocycle $\alpha_{\varepsilon}\in{\rm Z}^{2}_{\rm sym}(\mathcal{C}_{S},\{\pm 1\})$ by
@@
where $\mathop{\rm sgn}\nolimits$ denotes the usual sign function, which is the projection $\mathbb{R}^{\times}\twoheadrightarrow\{\pm 1\}$. The class $\alpha_{\varepsilon}]\in{\rm Ext}(\mathcal{C}_{S},\{\pm 1\})$ is the image of $\varepsilon]$ under the homomorphism ${\rm Hom}(\mathcal{I}_{S},\mathbb{R}^{\times})/{\rm B}^{2}_{S}(\mathbb{R}^{\times})\to{\rm Ext}(\mathcal{C}_{S},\{\pm 1\})$ from sequence (6). The corresponding element $\delta_{\varepsilon}\in{\rm Hom}((\mathcal{C}_{S})_{[2]},\{\pm 1\})$ is given by
@@ 4 Variety of graded contractions
Recall that Zariski topology on the affine space $\mathbb{F}^{n}$, $n\in\mathbb{N}$, is defined by taking as closed sets all algebraic sets, i.e., a set $V\subset\mathbb{F}^{n}$ is Zariski closed if it is the set of common zeros of a collection of polynomials from $R:=\mathbb{F}[x_{1},\dots,x_{n}]$. These polynomials generate an ideal $J$ of $R$, and $V$ is the vanishing set of $J$, $V(J):=\{x\in\mathbb{F}^{n}\mid f(x)=0\ \forall f\in I\}$. Conversely, for any set $S\subset\mathbb{F}^{n}$, we can consider its vanishing ideal, $I(S):=\{f\in R\mid f(x)=0\ \forall x\in S\}$. The set $V(I(S))$ is the Zariski closure of $S$, and the cornerstone result in affine algebraic geometry, Hilbert’s Nullstellensatz, says that, for any ideal $J$ of $R$, $I(V(J))=\sqrt{J}$, where the ideal $\sqrt{J}:=\{f\in R\mid\exists n\in\mathbb{N}\colon f^{n}\in J\}$ is called the radical of $J$. Thus we have a one-to-one correspondence between Zariski closed sets and radical ideals $I$ of the polynomial ring $R$, i.e., the ideals satisfying $I=\sqrt{I}$.
By definition, an affine algebraic variety $X$ is isomorphic to a Zariski closed set in the affine space $\mathbb{F}^{n}$ for some $n\in\mathbb{N}$, i.e., $X\simeq V(J)$ for some ideal $J\subset R$. The geometric properties of $X$ can be studied through the algebra of polynomial functions restricted to $X$, known as the algebra of regular functions and denoted $\mathbb{F}[X]:=R/\sqrt{J}$. For more details, see standard references like [8, 24].
@@
$x_{\{g,h\}}x_{\{gh,k\}}-x_{\{h,k\}}x_{\{g,hk\}}=0\quad\mbox{for all}\quad g,h,k\in G.$ (11)
Let $I\subset\mathbb{F}[\mathbb{A}_{G}]$ be the ideal generated by the left-hand sides of the above equations. This ideal is binomial, meaning it is generated by polynomials with two terms. Binomial ideals have a number of remarkable properties, for instance, the radical and the associated primes of a binomial ideal are again binomial. Eisenbud and Sturmfels [16] provide a detailed study of these and other properties and characterize algebraic sets defined by binomial ideals in terms of their cellular decomposition, see [16, Theorem 4.2]. We will subsequently study the cellular decomposition of the variety $X$ of graded contractions. It should be noted that the ideal $I$ is usually not radical, which can be verified using Singular computer algebra system [10]. For instance, $I$ is radical when $G$ is of order $2$, 33, 44 or $5$, but not $6$.
For any subset $S\subset\mathcal{P}_{G}$, we consider the affine subspace $\mathbb{A}_{S}:=\{x\in\mathbb{A}_{G}\colon x_{i}\neq 0~\forall i\in S\}\subset\mathbb{A}_{G}$ and the torus ${\rm T}_{S}:=\{x\in\mathbb{A}_{S}\colon x_{i}\neq 0~\forall i\in S\}$, which is a principal open set in $\mathbb{A}_{S}$. Since $x\in{\rm T}_{S}$ if and only if $\mathop{\rm supp}x=S$, where $\mathop{\rm supp}x=\{i\in\mathcal{P}_{G}\colon x_{i}\neq 0\}$, we have a partition $\mathbb{A}_{S}=\bigcup_{S^{\prime}\subset S}{\rm T}_{S^{\prime}}$. The tori ${\rm T}_{S}$ are called coordinate cells in [16]. The intersection $X\cap{\rm T}_{S}$ consists precisely of the graded contractions whose support is $S$; it is nonempty if and only if $S\in\mathcal{S}(G)$. It follows that we have the following partition of $X$:
@@
In particular, $X_{\mathcal{P}_{G}}=X\cap{\rm T}_{\mathcal{P}_{G}}={\rm Z}^{2}_{\rm sym}(G,\mathbb{F}^{\times})$ is a nonempty principal open set in $X$. Consider the torus ${\rm T}_{G}:={\rm Fun}(G,\mathbb{F}^{\times})$ and its action on the space $\mathbb{A}_{G}$ defined by
$\alpha\cdot(x_{\{g,h\}})_{\{g,h\}\in\mathcal{P}_{G}}=\big(d\alpha(g,h)\,x_{\{g,h\}}\big)_{\{g,h\}\in\mathcal{P}_{G}},$
where recall that $d\alpha(g,h)=\alpha(g)\alpha(h)/\alpha(gh)$. In other words, ${\rm T}_{G}$ acts through the homomorphism $d\colon{\rm T}_{G}\to{\rm T}_{\mathcal{P}_{G}}$, $\alpha\mapsto d\alpha$. This action leaves $X$ and each cell $X_{S}$ invariant; moreover, the group ${\rm H}^{2}_{S}(\mathbb{F}^{\times})$, which was introduced in the previous section, is in bijection with the set of orbits of $X_{S}$ under this action. Note that the ${\rm T}_{G}$-orbit of the point $\bar{1}:=(1,\dots,1)$ is ${\rm B}^{2}(G,\mathbb{F}^{\times})={\rm Z}^{2}_{\rm sym}(G,\mathbb{F}^{\times})=X_{\mathcal{P}(G)}$.
@@
The closure of the ${\rm T}_{G}$-orbit of $\bar{1}$, which we denote by $X(\bar{1})$, is the unique irreducible component of $X$ containing $\bar{1}$. Moreover, the dimension of $X(\bar{1})$ is $N=|G|$.
@@
Since ${\rm T}_{G}$ is irreducible, so is $X(\bar{1})$, hence it is contained in an irreducible component of $X$. But it contains an open neighborhood of $\bar{1}$ in $X$, so it must contain any irreducible component of $X$ passing through $\bar{1}$.
Since $\dim{\rm T}_{G}=N$ and the stabilizer of $\bar{1}$, ${\rm Hom}(G,\mathbb{F}^{\times})$, is finite, the second assertion follows from the fact that the ${\rm T}_{G}$-orbit of $\bar{1}$ is open in its closure, $X(\bar{1})$, see Section 2.2. ∎
@@
Consider the grading group $G=\mathbb{Z}_{2}$. The affine space $\mathbb{A}_{\mathbb{Z}_{2}}$ is coordinatized by the tuple $(x_{\{0,0\}},x_{\{1,1\}},x_{\{0,1\}})$ and its coordinate algebra is $\mathbb{C}[x_{\{0,0\}},x_{\{1,1\}},x_{\{0,1\}}]$. The variety of contractions is the zero locus of the polynomials
@@
The computations for the grading group $G=\mathbb{Z}_{3}$ are more complicated than those presented in Example 25. The coordinate algebra of $\mathbb{A}_{\mathbb{Z}_{3}}$ is the polynomial algebra in six variables, $\mathbb{C}[x_{\{0,0\}},x_{\{1,1\}},x_{\{2,2\}},x_{\{0,1\}},x_{\{0,2\}},x_{\{1,2\}}]$. The variety $X$ of contractions is the zero locus of the polynomials
@@
$x_{\{g,h\}}x_{\{gh,k\}}-x_{\{h,k\}}x_{\{g,hk\}},\quad x_{\{g,h\}}(x_{\{g,h\}}-1)\quad\mbox{for all}\quad g,h,k\in G.$
@@
All the above computations can be efficiently carried out in Singular. Even for the order-8 grading groups, i.e., the number of variables of $\mathbb{C}[\mathbb{A}_{G}]$ is $36$, these computations are performed within seconds. In Table 1 we present the numerical results for the number of supports depending on the grading group $G$. We also observe an interesting phenomenon that only a small part of subsets of $\mathcal{P}_{G}$ are actually supports.
@@
$2^{|\mathcal{P}(G)|}$ 8 $2^{6}$ $2^{10}$ $2^{10}$ $^{15} 22$$^{21}$ $2^{28}$ $2^{36}$ $2^{36}$ $2^{36}$
$\frac{|\mathcal{S}(G)|}{2^{|\mathcal{P}(G)|}}$ 0.625 $0.2344$ $0.040$ $0.046$ $1.43{\rm E}\mbox{-3}$ $3.11{\rm E}\mbox{-4}$ $1.12{\rm E}\mbox{-5}$ $2.35{\rm E}\mbox{-7}$ $2.06{\rm E}\mbox{-7}$ $2.02{\rm E}\mbox{-7}$
Orbit closures and higher-order identities. The elements of $X(\bar{1})$ are precisely the continuous contractions as defined in [36, Section 4]. In the case $\mathbb{F}=\mathbb{C}$, $\mathfrak{L}^{\varepsilon}$ is a graded degeneration of a Lie algebra $\mathfrak{L}$ for any $\varepsilon\in X(\bar{1})$, so, in particular, it is a contraction of $\mathfrak{L}$ in the sense of Definition 1. Recall from Section 3 that every relation $r=1$ for $r\in\langle\mathcal{R}_{G}\rangle$ can be uniquely rewritten in the form $r_{1}=r_{2}$, where $r_{1}$ and $r_{2}$ are monomials of the same degree in two disjoint sets of variables. Following [47, 48], we call these relations higher-order identities; they include the defining relations (11). Since the elements of ${\rm Z}^{2}_{\rm sym}(G,\mathbb{F}^{\times})$ satisfy all higher-order identities, so do the elements of the closure, $X(\bar{1})$. It follows from Theorem 19 and [47, Theorem VI.2] that any element of $X(\bar{1})$ is equivalent via normalization to a generalized Inönü–Wigner contraction.
@@
$\alpha\cdot(x_{\{g,h\}})_{\{g,h\}\in S}=\big(d\alpha(g,h)\,x_{\{g,h\}}\big)_{\{g,h\}\in S}.$
@@
The orbit of $\bar{1}_{S}$ is the range of the function $f_{S}\colon{\rm Fun}(G,\mathbb{F}^{\times})\to\mathbb{A}_{S}$, $\alpha\mapsto\alpha\cdot\bar{1}_{S}$. Since $\alpha\cdot\bar{1}_{S}=(d\alpha)\bar{1}_{S}$, the range of $f_{S}$ is precisely ${\rm B}^{2}_{S}(\mathbb{F}^{\times})$ defined in Section 3. For any monomial $r$ in variables indexed by $i\in S$, denote by $\lambda_{r}$ the restriction of the corresponding polynomial function $\mathbb{A}_{S}\to\mathbb{F}$ to the torus ${\rm Fun}(S,\mathbb{F}^{\times})$. First we claim that $\lambda_{r_{1}}|_{{\rm B}^{2}_{S}(\mathbb{F}^{\times})}=\lambda_{r_{2}}|_{{\rm B}^{2}_{S}(\mathbb{F}^{\times})}$ if and only if $r_{1}r_{2}^{-1}\in\mathcal{\langle}\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$.
@@
It is clear that $\lambda_{r_{1}}|_{{\rm B}^{2}_{S}(\mathbb{F}^{\times})}=\lambda_{r_{2}}|_{{\rm B}^{2}_{S}(\mathbb{F}^{\times})}$ if and only if the function $\lambda_{r_{1}}/\lambda_{r_{2}}$ maps ${\rm B}^{2}_{S}(\mathbb{F}^{\times})$ to $1$, and the latter condition is equivalent to $r_{1}r_{2}^{-1}\in{\rm B}^{2}_{S}(\mathbb{F}^{\times})^{\perp}\subset\mathcal{F}_{S}$.
Theorem 19 and Corollary 21 give us that the kernel of the restriction map ${\rm Hom}(\mathcal{Z}_{S},\mathbb{F}^{\times})\to{\rm Hom}(\mathcal{K}_{S},\mathbb{F}^{\times})$ is ${\rm B}^{2}_{S}(\mathbb{F}^{\times})$. Since $\mathcal{Z}_{S}=\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$ and $\mathcal{K}_{S}=(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})/\langle\mathcal{R}_{S}\rangle$, the kernel of the restriction map ${\rm Hom}(\mathcal{F}_{S},\mathbb{F}^{\times})\to{\rm Hom}(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S},\mathbb{F}^{\times})$ is also ${\rm B}^{2}_{S}(\mathbb{F}^{\times})$. By definition, this kernel is $(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})^{\perp}$, so we get ${\rm B}^{2}_{S}(\mathbb{F}^{\times})=(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})^{\perp}$ and hence ${\rm B}^{2}_{S}(\mathbb{F}^{\times})^{\perp}=((\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})^{\perp})^{\perp}$. Finally, $\mathcal{F}_{S}/(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})\simeq\mathcal{I}_{S}$ is a free abelian group, so its points are separated by homomorphisms to $\mathbb{F}^{\times}$. Therefore, $((\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})^{\perp})^{\perp}=\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$, which completes the proof of the claim.
Let $I$ be the vanishing ideal of the closure of ${\rm B}^{2}_{S}(\mathbb{F}^{\times})$ in $\mathbb{A}_{S}$. Since ${\rm B}^{2}_{S}(\mathbb{F}^{\times})$ is dense in its closure, we have
$I=\{\varphi\in\mathbb{F}[\mathbb{A}_{S}]\mid\varphi|_{{\rm B}^{2}_{S}(\mathbb{F}^{\times})}=0\}.$
A binomial $r_{1}-r_{2}$ arises from a surviving higher-order identity if and only if $r_{1}r_{2}^{-1}\in\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$. It follows from the above claim that all these binomials $r_{1}-r_{2}$ belong to $I$. Modulo a linear combination of such binomials, any $\varphi\in\mathbb{F}[\mathbb{A}_{S}]$ is congruent to a linear combination of monomials $r_{j}$ such that $r_{i}r_{j}^{-1}\notin\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$ for $i\neq j$. But the restrictions $\lambda_{r_{j}}|_{{\rm B}^{2}_{S}(\mathbb{F}^{\times})}$ are pairwise distinct multiplicative characters of ${\rm B}^{2}_{S}(\mathbb{F}^{\times})$, hence they are linearly independent over $\mathbb{F}$. Therefore the ideal $I$ is spanned by the binomials $r_{1}-r_{2}$ arising from surviving higher-order identities. This finishes the proof of the first claim of the theorem
@@
When $G$ is a finite group, the closure of the orbit of $\bar{1}_{S}$ is, by definition, the toric variety arising from the integral $|G|\times|S|$ matrix $A$, whose columns are indexed by $\{g,h\}\in S$ and are given by components of the vector g+hghg+h-gh in the standard basis of the free $\mathbb{Z}$-module $\mathbb{Z}G$, see [9, §1.1] for details about toric varieties. The nullspace $\{x\in\mathbb{Z}^{|S|}\mid Ax=0\}$ of the matrix $A$ can be identified with the abelian group $\mathcal{F}_{S}\cap\ker\partial_{2}$, where $\partial_{\bullet}$ is the differential of the complex $C_{\bullet}$ defined in Section 3. Since $\mathop{\rm im}\partial_{3}=\langle\mathcal{R}_{G}\rangle$ and $\mathcal{F}_{S}\cap\ker\partial_{2}=\mathcal{F}_{S}\cap\mathop{\rm im}\partial_{3}$ by exactness of the complex $C_{\bullet}$, Theorem 27 follows from [9, Proposition 1.1.9].
@@
The contracted algebra $\mathcal{L}^{\varepsilon}$ is a graded degeneration of $\mathcal{L}$ if and only if $\varepsilon$ satisfies all surviving higher-order identities for $S={\rm Supp}_{2}(\Gamma)$. In particular, if ${\rm Supp}_{2}(\Gamma)=\mathcal{P}_{G}$, then the graded degenerations are precisely the graded contractions in $X(\bar{1})$, i.e., they satisfy all higher-order identities.